# HypeLoRA: Hyper-Network-Generated LoRA Adapters for Calibrated Language Model Fine-Tuning

### 01 Mar 2026 v1

@@ 1 Introduction
Additionally, we study a modular calibration framework in which a lightweight hyper-network generates low-rank calibration signals conditioned on model structure. These signals are injected via low-rank adaptation modules, enabling context-dependent adjustment of confidence estimates while preserving the pretrained parameters of the base model. Although this approach ultimately fails to yield consistent calibration improvements, it provides valuable empirical insight into the limitations of hyper-network-driven calibration for transformer encoders.
@@
Overall, this work highlights both the promise and the limitations of parameter-efficient, input-dependent calibration mechanisms, and clarifies the trade-offs involved in moving beyond static post-hoc calibration for large transformer models.
@@ 3.2 Layer-Conditioned Low-Rank Updates with Hyper-Network
Unlike standard LoRA, which trains all of $A$ and $B$ matrices independently, we generate these factors via a shared hyper-network $H_{\phi}$ with parameters $\phi$. Each target weight matrix — specifically the Query and Value projections in each layer — is associated with a dedicated learned embedding $e\in\mathbb{R}^{d_{h}}$. Concretely, if both $A$s and $B$s are generated for the Query and Value matrices in layer $\ell$, there are four embeddings per layer: $e^{Q}_{A}[\ell],\,e^{Q}_{B}[\ell],\,e^{V}_{A}[\ell],\,e^{V}_{B}[\ell]$. Hyper-network $H_{\phi}$ maps each such embedding $e$ to target low-rank factor:
@@
  • Fixed-A: $A$ matrices are initialized once from a Kaiming uniform distribution (similarly to [10]) and held fixed throughout training; only $B$ matrices are generated by $H_{\phi}$, meaning that there are only two embeddings per layer.
@@ 5.2 Our Approach - Calibration with Hyper-Network
On CoLA, the Transformer-based hyper-network with fully generated matrices (Transformer $A_{\text{gen}}$) achieves average highest performance, but also shows high variability across different seeds. It surpasses LoRA fine-tuning. However, its calibration remains similar to LoRA, indicating no inherent calibration advantage from full matrix generation. When $A$ matrices are fixed (Transformer $A_{\text{fix}}$), MCC value decreases, but ECE improves, representing the best calibration among all evaluated configurations.

### 29 Mar 2026 v2

@@ 1 Introduction
Additionally, we study a framework in which a hyper-network generates low-rank calibration signals. This setup aims to induce structural coupling across layers, forcing the model to learn a global calibration strategy rather than independent adjustments. We anticipated that this shared generation process would act as a regularizer, tempering overly sharp probability distributions. While this approach ultimately fails to yield consistent improvements, it provides valuable insight into the limitations of hyper-network-driven calibration for transformers.
@@ 3.2 Layer-Conditioned Low-Rank Updates with Hyper-Network
Unlike standard LoRA, which trains all of $A$ and $B$ matrices independently, we generate these factors via a shared hyper-network $H_{\phi}$, conditioned on layer embedding. Each target weight matrix — specifically the Query and Value projections in each layer — is associated with a dedicated learned embedding $e\in\mathbb{R}^{d_{h}}$. Concretely, if both $A$s and $B$s are generated for the Query and Value matrices in layer $\ell$, there are four embeddings per layer: $e^{Q}_{A}[\ell],\,e^{Q}_{B}[\ell],\,e^{V}_{A}[\ell],\,e^{V}_{B}[\ell]$. Hyper-network $H_{\phi}$ maps each such embedding $e$ to target low-rank factor:
@@
  • Fixed-A: $A$ matrices are initialized once from a Kaiming uniform distribution (similarly to [10]) and held fixed throughout training; only $B$ matrices are generated by $H_{\phi}$, meaning that there are only two embeddings per layer. The intuition is that fixed random $A$ matrices act as structured noise injection into the frozen base model, encouraging the learned $B$ matrices to be more robust.
@@ 5.2 Our Approach - Calibration with Hyper-Network
On CoLA, the Transformer-based hyper-network with fully generated matrices (Transformer $A_{\text{gen}}$) achieves average highest performance, but also shows high variability across different seeds. It surpasses LoRA fine-tuning. However, its calibration remains similar to LoRA, indicating no inherent calibration advantage from full matrix generation. When $A$ matrices are fixed (Transformer $A_{\text{fix}}$), MCC value decreases, but ECE improves, representing the best calibration among all evaluated configurations. We hypothesize that this calibration improvement is caused by a reduction in effective degrees of freedom - fixing $A$ matrices constrains the adaptation space, acting as an implicit regularizer that prevents the model from becoming overconfident.